### v2

# Maximal work extraction unitarily from an unknown quantum state: Ergotropy estimation via feedback experiments

Abstract
Considering the emerging applications of quantum technologies, studying energy storage and usage at the quantum level is of great interest. In this context, there is a significant contemporary interest in studying ergotropy, the maximum amount of work that can be extracted unitarily from an energy-storing quantum device. Here, we propose and experimentally demonstrate a feedback-based algorithm (FQErgo) for estimating ergotropy. This method also transforms an arbitrary initial state to its passive state, which allows no further unitary work extraction. FQErgo applies drive fields whose strengths are iteratively adjusted via certain expectation values, conveniently read using a single probe qubit. Thus, FQErgo provides a practical way for unitary energy extraction and for preparing passive states. By numerically analyzing FQErgo on random initial states, we confirm the successful preparation of passive states and estimation of ergotropy, even in the presence of drive errors. Finally, we implement FQErgo on two- and three-qubit NMR registers, prepare their passive states, and accurately estimate their ergotropy.
I Introduction
Quantum technologies are appealing as they can surpass their classical counterparts by using exclusive concepts like coherence and entanglement. Understanding their capabilities has progressed in recent years, thanks to the numerous experimental breakthroughs that have improved control over quantum states. Specifically, the investigation of the principles of quantum thermodynamics is made possible by quantum thermal machines, which include refrigerators and heat engines to regulate heat flow and work production [1, 2]. Further options for storing energy for later extraction include quantum batteries [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. The quantum batteries have the potential to outperform their classical counterparts in relevant metrics like charging speed [14, 15, 16, 17, 18, 19, 20], stored energy [21, 22, 20], and energy extraction [23, 24, 25, 26, 27, 28, 29].
The maximum amount of energy that can be extracted from quantum systems through unitary processes is known as the ergotropy [30]. A practical way to determine ergotropy is by finding the optimal unitary operation that transforms the system to its lowest attainable energy state, known as its passive state [31]. This can be a challenging undertaking as ergotropy might be sensitive to correlations that can also impact device performance [32, 33, 34, 35, 36, 37, 38, 39]. Recent works have tried to establish a link between entanglement and thermodynamic quantities like the ergotropy gap, the difference between the maximal extractable works via global and local unitaries. These thermodynamic quantities can be used to verify entanglement present in bipartite [40], multipartite [41, 42], and multiqubit mixed states [43]. Using these quantities, experimental certification of entanglement has been demonstrated in up to 10 qubits [43].
The main focus of this work is FQErgo, an algorithm we propose for preparing the passive state and thereby estimating ergotropy. FQErgo is inspired by (hence the name) the feedback-based algorithm for quantum optimization (FALQON) [44], which has been recently developed for combinatorial quantum optimization. This method is related to quantum Lyapunov control (QLC) and uses feedback conditioned on qubit measurements at each quantum circuit layer and parametrized quantum circuits to determine the values of the circuit parameters at subsequent layers [45, 46].
We note several salient features of the FQErgo algorithm. Firstly, the all-quantum nature of the algorithm: unlike the previous methods [47, 48] for estimating ergotropy, FQErgo requires no classical optimizer, and can be fully automated for convenient practical realizations. Secondly, FQErgo is robust against circuit errors since each feedback iteration readjusts the drives based on previous errors. This way, cumulative error growth is prevented. Thirdly, the circuit parameters can be read efficiently via an ancillary probe qubit using the interferometric method [49, 50, 51], which alleviates the need for resource-intensive quantum state tomography. In this work, after numerically analyzing the FQErgo algorithm, we experimentally demonstrate it using two and three-qubit NMR registers.
This article is organized as follows. In Sec. II, we describe the theoretical construction of FQErgo algorithm. In Sec. III, we numerically analyze FQErgo over random sets of initial states in one and two-qubit systems. The NMR implementations of FQErgo for both one and two-qubit systems with an additional probe qubit are reported in Sec. IV. Finally, Sec. V summarizes the results and discusses their importance.
II Theory
II.1 Unitary extraction of work, passive state, and ergotropy
Given a quantum system in a state $\rho_{0}$, the ergotropy quantifies the maximum amount of work that is unitarily extracted while leaving the system into a passive state $\rho^{p}$ from which no further work can be unitarily extracted [30]. If $H_{0}$ is the system Hamiltonian, the ergotropy is given by
${\cal E}(\rho_{0})$ $=E(\rho_{0})-E(\rho^{p}),~{}\mbox{where,}$,
$E(\rho^{p})$ $=\min_{U}\mathrm{Tr}\left(U\rho_{0}U^{\dagger}H_{0}\right).$ (1)
Let ${\cal U}_{p}$ corresponds to the unitary that minimizes the system energy such that $\rho^{p}={\cal U}_{p}\rho_{0}{\cal U}_{p}^{\dagger}$ is the passive state [30, 31]. If $H_{0}$ and $\rho_{0}$, in their respective eigenbases, are in the form
$H_{0}$ $=\sum_{k}E_{k}|{E_{k}}\rangle\langle{E_{k}}|,~{}\mbox{with}~{}E_{1}\leq E_{2}\leq\cdots~{}\mbox{and}$
$\rho_{0}$ $=\sum_{j}r_{j}|{r_{j}}\rangle\langle{r_{j}}|,~{}\mbox{with}~{}r_{1}\geq r_{2}\geq\cdots,$ (2)
then the passive state is uniquely defined in the energy eigenbasis as
$\rho^{p}$ $=\sum_{j}r_{j}|{E_{j}}\rangle\langle{E_{j}}|.$ (3)
While the above definitions of the passive state and ergotropy are simple and clear, the experimental realization of the passive state from an unknown initial state and thereby estimating its ergotropy has not been reported. In the following, we describe a robust method for the same.
II.2 FQErgo: A feedback algorithm for ergotropy estimation
Figure 1: (a) Illustrating $n$ iterations of FQErgo applied on an initial state $\rho_{0}$ to reach $\rho_{n}\approx\rho^{p}$, its passive state. (b,c) Determining FQErgo drive amplitude $\beta$ with a probe qubit, for (b) unitary Hermitian drive $\sigma_{\gamma}$, and (c) nonunitary Hermitian drive $A$. For energy measurement, we set $A=H_{0}$.
FQErgo is based on the FALQON algorithm proposed recently [44]. We assume multiple copies of individually accessible systems each prepared in the state $\rho_{0}$, which is not necessarily known. Alternatively, we can assume an oracle that transforms a uniquely initialized quantum system into a definite unknown state $\rho_{0}$. On introducing a drive Hamiltonian $\beta(t)H_{d}$, the system evolves as
$i\frac{d}{dt}\rho(t)=\left[H_{0}+\beta(t)H_{d}\,,\,\rho(t)\right],$ (4)
where $\beta(t)$ is the time-dependent drive parameter. Here and in the rest of the paper we have set $\hbar=1$. We seek to minimize the system energy $E(\rho(t))=\left\langle{H_{0}}\right\rangle_{\rho(t)}=\mathrm{Tr}(H_{0}\rho(t))$, which can be accomplished by designing control $\beta(t)$ satisfying the Lyapunov condition
$\frac{d}{dt}\left\langle{H_{0}}\right\rangle_{\rho(t)}=\beta(t)\,\left\langle{C_{d}}\right\rangle_{\rho(t)}\leq 0,~{}~{}\forall~{}t\geq 0$ (5)
where $C_{d}=i[H_{d},H_{0}]$. One way to satisfy the above is by choosing the control in the form [44]
$\beta(t)=-w\left\langle{C_{d}}\right\rangle_{\rho(t)},$ (6)
where $w$ is a positive scalar coefficient.
The overall algorithm is illustrated in Fig. 1 (a). We apply discrete drives each of duration $\tau$ and realize a $n$-step feedback loop,
$U_{n}=u_{n}\cdots u_{2}u_{1},~{}\mbox{with}~{}~{}u_{k}=e^{-i\beta_{k}\tau H_{d}}.$ (7)
We find the first drive parameter by measuring the expectation value $\beta_{1}=-w\,\left\langle{C_{d}}\right\rangle_{\rho_{0}}$, and prepare the state $\rho_{1}=U_{1}\rho_{0}U_{1}^{\dagger}$. The subsequent steps involve finding
$\beta_{k}=-w\,\left\langle{C_{d}}\right\rangle_{\rho_{k-1}}~{}\mbox{and preparing}~{}\rho_{k}=U_{k}\rho_{0}U_{k}^{\dagger}.$ (8)
Thus, starting from an unknown quantum state $\rho_{0}$, using the sequence of feedback-designed operators $U_{n}\approx{\cal U}_{p}$, we attain the minimum saturated energy state $\rho_{n}\approx\rho^{p}$. The energy measurements of the system for the initial state $\rho_{0}$ and the final state $\rho_{n}\sim\rho_{p}$ yield an estimation of ergotropy via Eq. 1.
II.3 FQErgo using a probe qubit
[…]
$\left\langle{\sigma_{x}}\right\rangle_{\mathrm{probe}}=\left\langle{U}\right\rangle_{\rho}.$ (9)
Thus, we directly obtain the desired expectation value of the system as the signal $\left\langle{\sigma_{x}}\right\rangle_{\mathrm{probe}}$ in the probe qubit. For example, if the system Hamiltonian is the Pauli operator $\sigma_{z}$ as described later in the experimental section, the probe signal directly yields the expectation value $\left\langle{\sigma_{z}}\right\rangle_{\rho}$. For the more general case of extracting the expectation value $\left\langle{U}\right\rangle_{\rho}$ of a nonunitary Hermitian observable $A$, we construct a unitary $e^{-i\alpha A}$, where $\alpha$ is a real parameter such that $||\alpha A||\ll 1$. Now implementing the controlled gate as shown in Fig. 1 (c), we obtain the probe signal [50]
$\left\langle{\sigma_{y}}\right\rangle_{\mathrm{probe}}=-\left\langle{\sin(A\alpha)}\right\rangle_{\rho}\approx-\alpha\left\langle{A}\right\rangle_{\rho}.$ (10)
Here, by setting $A=C_{d}$, we can extract $\left\langle{C_{d}}\right\rangle_{\rho}$ directly as the probe signal $\left\langle{\sigma_{y}}\right\rangle_{\mathrm{probe}}$.
Figure 2: (a,b) Simulated system-energy $E(\rho_{k})$ vs iteration number $k$ for a single qubit ($\omega_{0}=1)$ starting from 20 random initial states applied with (a) ideal FQErgo and (b) FQErgo having random errors. Insets show the first three iterations. (c) Numerically estimated vs exact ergotropy values with ideal FQErgo (filled circles) and FQErgo with random errors (filled triangles).
III Numerical simulations
III.1 Single qubit system
We consider a single qubit system with Hamiltonian $H_{0}=\omega_{0}(\mathrm{I}-\sigma_{z})/2$ with energy eigenvalues $(0,\omega_{0})$ prepared in an arbitrary mixed state $\rho(0)=(1-\epsilon)/2\mathrm{I}+\epsilon|{\psi}\rangle\langle{\psi}|$, where $|{\psi}\rangle=\cos(\theta/2)|{0}\rangle+e^{i\phi}\sin(\theta/2)|{1}\rangle$ and the purity parameter $\epsilon\in[0,1]$. We now use FQErgo to reach its passive state $\rho^{p}=(1-\epsilon)\mathrm{I}/2+\epsilon|{0}\rangle\langle{0}|$. For the one qubit case, FQErgo needs only two drives: $\sigma_{x}$ and $\sigma_{y}$. Fig. 2 (a) shows the energy $E(\rho_{k})$ versus iteration number $k$ for a set of 20 random initial states. In all cases, we see monotonically decreasing energy, ultimately approaching their respective passive states for sufficiently large iterations $n$, i.e., $\rho_{n}\rightarrow\rho^{p}$. The difference between the initial energy $E(\rho_{0})$ and the final energy $E(\rho_{n})$ estimates the ergotropy, i.e., ${\cal E}_{\mathrm{est}}(\rho_{0})=E(\rho_{0})-E(\rho_{n})$. The filled circles in Fig. 2 (c), plotting ${\cal E}_{\mathrm{est}}(\rho_{0})$ vs ${\cal E}_{\mathrm{exact}}(\rho_{0})=E(\rho_{0})-E(\rho^{p})$ confirm perfect estimation of ergotropy for all the random initial states.
Now we numerically analyze the robustness of FQErgo against potential imperfections in practical implementations. Firstly, being a feedback-based method, FQErgo is naturally robust against the drive amplitude. If the drive amplitude is lesser than the nominal, it still reaches the passive state but simply takes more iterations. To study the effect of more general errors, we introduce an error rotation of 5 degrees about a random direction in every FQErgo iteration. The system-energy curves in Fig. 2 (b) no longer exhibit a perfect monotonic decay, but neither show serious build-up of errors. The corresponding ergotropy values shown by filled triangles in Fig. 2 (c) show somewhat underestimated values, which is expected since circuit errors can only make the system settle in a higher energy state by preventing it from reaching the passive state.
Figure 3: (a,b) FQErgo circuits for local (a) and global (b) energy extraction in two-qubits. Here, local drives $X$, $Y$ suffice for reaching the local passive state $\rho^{(12)L}$ and estimating the local ergotropy ${\cal E}^{L}(\rho^{(12)}_{0})$ (a). In contrast, additional global drive such as $XY$ and $Z_{\delta}$ [47] gates are necessary for reaching the global passive state $\rho^{(12)G}$ and estimating the global ergotropy ${\cal E}^{G}(\rho^{(12)}_{0})$ (b). (c,d) Simulated system-energy vs iteration number for the two-qubit ($\omega_{0}=1$ and $J=0.01$) FQErgo from 20 random initial states without error (c) and with error (d). (e,f) The numerically estimated local (e) and global ergotropy (f) vs the exact values. (g) Ergotropy gap obtained from (e) and (f), vs the exact values. In (e-g), the filled circles are without random errors and filled triangles are with random errors.
III.2 Two qubit system
There is a fundamental difference between work extractions from one-qubit and multi-qubit systems. For two or more qubits, energy can be extracted using local unitaries on individual systems or global unitaries on the whole system. If we minimize the system energy using only local unitaries, we can get to the local passive state $\rho^{p_{L}}$, which determines the local ergotropy ${\cal E}^{L}=E(\rho_{0})-E(\rho^{p_{L}})$. If the minimization controls involve global unitaries, in addition to local, we can get to the global passive state $\rho^{p_{G}}$, which determines the global ergotropy ${\cal E}^{G}=E(\rho_{0})-E(\rho^{p_{G}})\leq E(\rho^{p_{L}})$. The difference $\Delta={\cal E}^{G}-{\cal E}^{L}$ is called the ergotropy gap [52].
Consider a two-qubit system with Hamiltonian $H_{0}=\omega_{0}(\mathrm{I}\otimes\mathrm{I}-(\sigma_{z}\otimes\mathrm{I}+\mathrm{I}\otimes\sigma_{z})/2)+J(\sigma_{z}\otimes\sigma_{z})$ with $|J|\ll|\omega_{0}|$ and prepared in random initial states $\rho_{0}^{(12)}$. The ergotropy gap is now
${\Delta={\cal E}^{G}(\rho^{(12)}_{0})-({\cal E}^{L}(\rho^{(1)}_{0})+{\cal E}^{L}(\rho^{(2)}_{0}))}.$ (11)
Our goal is to use FQErgo for estimating both local and global ergotropies and thereby obtaining the ergotropy gap. As illustrated in Fig. 3 (a), we use four local drives $H_{\gamma}^{(1)}=\sigma_{\gamma}\otimes\mathrm{I}$ and $H_{\gamma}^{(2)}=\mathrm{I}\otimes\sigma_{\gamma}$, where $\gamma\in[x,y]$ over the first 30 iterations that saturated the energy and reached $\rho_{30}\approx\rho^{(12)p_{L}}$, the local passive state. From the 31st iteration, we see a further reduction in energy as we apply a combination of global and local operations, as illustrated in Fig. 3 (b). It involves a tilted phase gate
$Z_{\delta}=e^{-i\delta(I\otimes\sigma_{y})}\left[|{0}\rangle\langle{0}|\otimes I+|{1}\rangle\langle{1}|\otimes e^{-i\sigma_{z}\pi/2}\right]e^{i\delta(I\otimes\sigma_{y})}$
with a fixed small angle $\delta$ [47], local gates, as well as the global $XY$ gate,
$U_{xy}=e^{-i\beta^{xy}H_{xy}\tau_{xy}}~{}\mbox{with}~{}H_{xy}=\sigma_{x}\otimes\sigma_{y}+\sigma_{y}\otimes\sigma_{x},$
strength $\beta^{xy}$, and duration $\tau_{xy}$. Finally, we find the second energy saturation over 60 iterations while reaching $\rho_{60}\approx\rho^{(12)p_{G}}$, the global passive state. Fig. 3 (c) shows the monotonic energy decrease for 20 random initial states undergoing FQErgo iteration. For each initial state, we observe two minimum-energy states, one corresponding to the local passive state and the other to the global passive state. Figs. 3 (e-f) plot the estimated local and global ergotropies against their exact values, and Fig. 3 (g) plots the estimated ergotropy gaps against exact values. Indeed, one can apply global energy extraction sequence Fig. 3 (b) directly from initial point and reach passive state much faster but here our focus is to estimate the ergotropy gap $\Delta$.
We again study FQErgo robustness by introducing a random nonlocal error unitary $U_{err}=e^{-iH_{err}\eta}$ generated by a random unit-norm error Hamiltonian $H_{err}$ and $\eta\equiv 2$ degrees in every FQErgo iteration. The system-energy curves shown in Fig. 3 (d) lose monotonicity in decay, but retain the overall trend. The error-affected local and global ergotropy values and the corresponding ergotropy gaps are shown by filled triangles in Fig. 3 (e-g). Likewise in one qubit case, the ergotropy values, particularly the global ergotropy values are underestimated, but the rms deviation of the ergotropy gap estimations from the ideal values remains below 0.07 indicating reasonable robustness of FQErgo against unitary errors.
IV Experiments
IV.1 Single-qubit ergotropy
We now use sodium fluorophosphate (dissolved in D2O; Fig. 4(a)) as the two-qubit register, wherein 19F nuclear spin is our probe qubit and 31P nuclear spin is the system qubit. All the experiments were carried out on a $500$ MHz Bruker NMR spectrometer at an ambient temperature of $300$ K. The rotating frame Hamiltonian of the system consisting of the internal part and the RF drive is $H^{\mathrm{FP}}=H^{\mathrm{FP}}_{\mathrm{int}}+H^{\mathrm{FP}}_{\mathrm{RF}}(t)$, where $H^{\mathrm{FP}}_{\mathrm{int}}=-\omega_{\mathrm{F}}I_{z}^{\mathrm{F}}-\omega_{\mathrm{P}}I_{z}^{\mathrm{P}}+2\pi J_{\mathrm{FP}}I_{z}^{\mathrm{F}}I_{z}^{\mathrm{P}}$ and $H^{\mathrm{FP}}_{\mathrm{RF}}(t)=\Omega_{\mathrm{F}}(t)I_{x}^{\mathrm{F}}+\Omega_{\mathrm{P}}(t)I_{x}^{\mathrm{P}}$, where $I_{x,y,z}$ are the spin operators. Here $(\omega_{\mathrm{F}}$, $\omega_{\mathrm{P}})$ and ($\Omega_{\mathrm{F}}(t)$, $\Omega_{\mathrm{P}}(t)$) respectively denote the adjustable frequency offsets and time-dependent RF amplitudes of 19F and 31P spins, and $J_{\mathrm{FP}}=868.0$ Hz is the scalar coupling constant.
Figure 4: (a) Sodium fluorophosphate molecular structure with $J_{\mathrm{FP}}$ indicated. (b) The experimentally measured one-qubit energy (normalized) vs FQErgo iterations for three initial states $\rho_{\theta,\phi}=|{0}\rangle\langle{0}|\otimes U_{\theta,\phi}|{0}\rangle\langle{0}|U^{\dagger}_{\theta,\phi}$ prepared by one of the RF rotations $U_{\theta,\phi}=e^{-i\theta(I_{x}\cos\phi+I_{y}\sin\phi)}\equiv(\theta,\phi)$ as indicated. (c) Ergotropy estimated from (b) vs simulated values. (d) Decaying fidelities of initial states ($F_{0}(\rho_{k})$, open symbols) and corresponding growth of passive states ($F_{p}(\rho_{k})$, filled symbols) for the same three cases as in (b).
The pulse sequences for the NMR implementations of FQErgo are described in Appendix. Starting from the thermal state, we prepare a pseudopure state (PPS) before energizing the system qubit 31P into one of the three initial states $\rho_{0}$ as described in Fig. 4. For FQErgo, we randomly choose the initial drives $\beta^{x}_{0}I_{x}$ and $\beta^{y}_{0}I_{y}$, and their strengths in the $k$th iteration are obtained by measuring the ancilla signal $\left\langle{i[H_{0},I_{x(y)}]}\right\rangle_{\rho_{k-1}}=\beta^{x(y)}_{k}$. Notice that all three different states reach their common passive state within 10 iterations in Fig. 4 (b). The corresponding ergotropy values show excellent agreement with the numerically simulated values in Fig. 4 (c), thus demonstrating the successful implementation of FQErgo.
In this one qubit case, since we have now measured all three expectation values $\left\langle{\sigma_{x,y,z}}\right\rangle_{\rho_{k}}$, we can now reconstruct the density matrix and determine its state fidelities $F_{0}(\rho_{k})=\mathrm{Tr}(\rho_{0}\rho_{k})$ with the respective initial state, and $F_{p}(\rho_{k})=\mathrm{Tr}(\rho_{p}\rho_{k})$ with the expected target passive state, at each iteration. The resulting fidelity profiles are shown in Fig. Fig. 4 (d). They indicate the gradual decay of the initial state fidelity and the corresponding buildup of the passive state fidelity. We clearly find high-fidelity passive states being prepared in each of the three cases.
IV.2 Two-qubit ergotropy
To estimate two-qubit ergotropy, we use the three-spin NMR register dibromofluoromethane (dissolved in Acetone-D6; Fig. 5 (a)). We consider 13C as the probe and 1H, 19F as the system qubits. The rotating frame Hamiltonian consisting of the internal part and the RF drive reads as $H^{\mathrm{CHF}}=H^{\mathrm{CHF}}_{\mathrm{int}}+H^{\mathrm{CHF}}_{\mathrm{RF}}(t)$, where $H^{\mathrm{CHF}}_{\mathrm{int}}=-\omega_{\mathrm{C}}I_{z}^{\mathrm{C}}-\omega_{\mathrm{H}}I_{z}^{\mathrm{H}}-\omega_{\mathrm{F}}I_{z}^{\mathrm{F}}+2\pi J_{\mathrm{CH}}I_{z}^{\mathrm{C}}I_{z}^{\mathrm{H}}+2\pi J_{\mathrm{CF}}I_{z}^{\mathrm{C}}I_{z}^{\mathrm{F}}+2\pi J_{\mathrm{HF}}I_{z}^{\mathrm{H}}I_{z}^{\mathrm{F}}$ and $H^{\mathrm{CHF}}_{\mathrm{RF}}(t)=\Omega_{\mathrm{C}}(t)I_{x}^{\mathrm{C}}+\Omega_{\mathrm{H}}(t)I_{x}^{\mathrm{H}}+\Omega_{\mathrm{F}}(t)I_{x}^{\mathrm{F}}$. Here $(\omega_{\mathrm{C}}$, $\omega_{\mathrm{H}},\omega_{\mathrm{F}})$ and $(\Omega_{\mathrm{C}},\Omega_{\mathrm{H}},\Omega_{\mathrm{F}})$, respectively denote the frequency offsets and RF amplitudes, while $(J_{\mathrm{CH}},J_{\mathrm{CF}},J_{\mathrm{HF}})$ denote the scalar coupling constants, whose values are shown in Fig. 5 (a).
Figure 5: (a) Molecular structure of dibromofluoromethane, with various $J$ couplings indicated. (b) The experimentally measured two-qubit energy (normalized) vs FQErgo iterations for five initial states with varying entanglement entropy ($S$) as described in the text. (c,d) Local (c) and global (d) ergotropy experimentally estimated from (b) vs simulated values. (d) The experimental ergotropy gap obtained from (c,d) vs simulated values.
To demonstrate local and global ergotropy estimation, we prepare five initial states with varying degrees of entanglement between the system qubits $\mathrm{H}$ and $\mathrm{F}$. The full NMR pulse-sequences are described in the Appendix. After preparing $|{000}\rangle\langle{000}|$ we transform it to $|{0}\rangle\langle{0}|\otimes\rho^{\mathrm{HF}}_{0}=|{0}\rangle\langle{0}|\otimes U_{G}|{00}\rangle\langle{00}|U_{G}^{\dagger}$ using a global unitary $U_{G}$ consisting of a Hadamard operator on $\mathrm{H}$ and a controlled $e^{-i\nu I_{x}}$ gate on $\mathrm{F}$, where $\nu$ controls the degree of entanglement. That completes the initialization, and now we start work extraction by FQErgo. During the first 10 iterations, we extract work by using only local drives $I_{x(y)}^{\mathrm{H}}$, $I_{x(y)}^{\mathrm{F}}$ (similar as in Fig. 3 (a)) and realize the first energy minimization corresponding to the local passive state as shown in Fig. 5 (c). From 11th to 20th iterations, we extract further energy via the global drive $I_{x}^{\mathrm{H}}I_{y}^{\mathrm{F}}+I_{y}^{\mathrm{H}}I_{x}^{\mathrm{F}}$ along with the $Z_{\delta}$ gate and local drives as shown in Fig. 3 (b). The global drive allows complete work extraction, eventually taking the system qubits to the second energy minimization corresponding to the global passive state $\rho^{p}_{\mathrm{HF}}$. Fig. 5 (b) shows the normalized energies for both local and global extraction with initial states having varying entanglement entropy $S=-\mathrm{Tr}[\rho^{\mathrm{H}}_{0}\log\rho^{\mathrm{H}}_{0}]$, where $\rho^{\mathrm{H}}_{0}=\mathrm{Tr}_{\mathrm{F}}(\rho_{0}^{\mathrm{HF}})$. Fig. 5 (c-d) plot the experimentally estimated local and global ergotropy values vs simulated values and Fig. 5 (e) plots the experimentally obtained ergotropy gaps vs versus the simulated values. The good agreement between the estimated values and the simulated values confirms the successful demonstration of the FQErgo algorithm.
V Summary and discussions
In summary, we have introduced a feedback-based algorithm, FQErgo to prepare the passive state and thereby quantify the ergotropy, the maximum unitarily extractable energy, of an unknown quantum state of a given system. It is an iterative feedback algorithm that efficiently reads certain expectation values using a probe qubit and readjusts the subsequent drive strengths to extract further energy. The same probe qubit also allows regular monitoring of system energy throughout the process. By numerically implementing FQErgo over a set of random states in both one and two-qubit systems, we have verified robust passive state preparation and ergotropy estimation, even in the presence of circuit errors. We then experimentally implemented FQErgo on multiple initial states of two and three-qubit NMR registers, successfully prepared their passive states, and closely estimated their local and global ergotropies as well as the ergotropy gap.
We envisage several future directions. For instance, while ergotropy based thermodynamic quantities can form convenient entanglement witnesses, they required prior knowledge of the class of states (eg. [43]). FQErgo can overcome such limitations and pave the way to certify the entanglement for a completely unknown state. Since one probe-qubit suffices, irrespective of the system size, extending FQErgo to larger systems should be feasible without exponential complexity. The procedure for extracting work from an unknown state is crucial from the perspective of quantum batteries. Although a fully charged quantum battery may start from a known state, subsequently during storage or partial usage it may end up in an unknown state. Although the ensemble nature of NMR is advantageous here, the overall algorithm is general enough to adapt to other quantum architectures.
Acknowledgements
Authors acknowledges valuable discussions with Pranav Chandarana, Mir Alimuddin, Vishal Varma, Arijit Chatterjee, Harikrishnan, and Conan Alexander. J.J. acknowledges support from CSIR (Council of Scientific and Industrial Research) Project No. 09/936(0259)/2019 EMR - I. We also thank the National Mission on In- terdisciplinary Cyber-Physical Systems for funding from the DST, Government of India, through the I-HUB Quantum Technology Foundation, IISER-Pune and funding from DST/ICPS/QuST/ 2019/Q67.
Appendix: NMR pulse sequences
Figure 6: NMR pulse sequence for the 1-qubit FQErgo experiment. From thermal equilibrium state, we first prepare the $|{00}\rangle$ pseudopure state (PPS) which is rotated by $U_{\theta,\phi}=e^{-i\theta(I_{x}\cos\phi+I_{y}\sin\phi)}$ to realize the initial state $\rho_{0}$. After FQErgo iterations $u_{k}=e^{-i\beta^{x}_{k}I_{x}\tau}e^{-i\beta^{y}_{k}I_{y}\tau}$ on the system qubit 31P (only one iteration is shown) and Hadamard gate (H) on the probe, we extract (i) one of the drive amplitudes $\beta^{x(y)}$ or (ii) energy $E$. Here the measurement pulse-sequences implement one of $e^{-i\beta^{x(y)}I_{y(x)}\tau}$ and $e^{-iI_{z}\tau}$ acting on system 31P qubit and controlled by 19F qubit. Finally, $\sigma_{x}$ measurements of 19F qubit yields drive amplitudes $\beta^{x(y)}$ or energy $E$.
Figure 7: The NMR pulse sequences for 2-qubit (1H,19F), FQEergo with one (13C) probe qubit: (a) PPS preparation, initializing the system as $\rho^{\mathrm{HF}}_{0}$ using entangling unitary $U_{G}$ after a pseudo-Hadamard on 1H. Here delay $d\tau$ is varied to change entanglement entropy. (b) Local and global drives, where $\delta_{y}=e^{-i\delta I^{F}_{y}}$, $U_{xy}=e^{-i\beta^{xy}H_{xy}\tau_{xy}}$ and the delay $d\beta$ controls the strength of the global drive. (c) Extracting global drive strength $\beta^{xy}$ and system energy.